A sample of hydrogen gas at a pressure of 733 mm Hg and a temperature of 25°C, occupies a volume of 13.7 liters. If the gas is heated at constant pressure to a temperature of 49°C, the volume of the gas sample will be


1
Expert's answer
2021-07-12T02:59:27-0400

PV=nRT

P - pressure

V - volume of gas

n - number if moles of gas

R-universal gas constant

T - temperature

P, n, R - are constants in our task.

PV1=nRT1

P/(nR)=T1/V1=(25+273)/13.7=21.75 K/L=const

25+273 is convertation from degrees to kelvins.

V2=T2/const=(49+273)/21.75=61.6 L